OpenStudy (anonymous):

I don't understand how to solve this at all, is there anybody who can explain? Please? Suppose 0.09886 M KOH is titrated into 15.00 mL H2SO4 of unknown concentration until the equivalence point is reached. It takes a total volume of 29.25 mL KOH to neutralize the H2SO4. How many moles of acid were neutralized during the titration? Express the answer to five significant figures.

OpenStudy (anonymous):

When it comes to solutions, use the formula n=cv/1000, where n is number of moles, v is volume in cm^3=mL. First write the balnced equation and balance it, so that you will know the mol ratio of the substances involved: 2KOH + H2SO4 - K2SO4 + 2H2O For first calculation, you should use the data of the substance that it is in limiting amount and in this case is KOH. So using the formula from above, we find its moles= 0.09886*29.25/1000=2.8917 mol

OpenStudy (anonymous):

Back to your balanced equation, you can see that KOh and H2SO4 are in the ratio of 2:1, hence the moles of the acid is half of those of KOH.

OpenStudy (anonymous):

Thank you that really helped.